OpenStudy (anonymous):

Solve the system algebraically. 2x + y - 10 = 0 x - y - 5 = 0 What is the value of y? -5/3 0 5

OpenStudy (anonymous):

i am having trouble with this

OpenStudy (anonymous):

First, make sure it is set up properly. You want to move the -10 and the 10 to the other side. 2x + y=10 x - y = 5 You want to eliminate the x. To do that, multiply the bottom equation by -2. You should get, (-2x+2y=-10) Now, you cancel out the x's, add the y's and the solutions. 2x+y=10 -2x+y=-10 You should end up with, 3y=0 Now divide by 3 to get the answer.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

No problem :)